------------------------------------------------------------ Date: Mon Aug 28 14:23:00 2000, Shift: 08:00-16:00 Supervisor: JJG, with: Log book#5 pages 36 - 37 Shift Summary: ------------------------------------------------------------ 1. Vaccum problems in Yellow ring cause large backgrounds. Y-Intensity:13 10**9, B-intensity: 17**9. ZDC singles rate is 96k/s in Y, 75k/s in B. With RHIC in coinc. rates are Y: 1.3k, B: 3.4k. Rate of trigger 6 is 109k/sec. At these rates a good part of the data may be junk. How bad it is needs to be evaluated. Decision is made not to run until vacuum improves (10.40). This vaccum problem mainly affects BRAHMS. 2. Only one short run taken. Run 2458 with 83kevents from 10.20 to 10.40. 3. Under above conditions TPC leak currents are: MTPC1: 500nA, MTPC2: 2000nA, T1: 1400nA, T2A: 3900nA, T2B:1300nA, i.e. 50-100 percent above 'normal'. Current magnets settings are MRS 90 deg. pol B, FS 5 deg pol B. Magnets are left on Stdby. TPC anodes and field wires are off. I have changed the trigger scale down factors. I cannot find a document listing the 'recommended' or 'standard' values that minimize DT. Summary of data accumulated so far. MRS 90 deg pol A 2kG : 5.0M trig 1 MRS 90 B : 3.8M trig 1 MRS 45 deg A : 1.5M trig 1 During this FS has been mostly at 5 deg at 1/5 full field. Estimate that we need another 6-8 days of good running to achieve goal of 10M MB events at 90 deg at each polarity (estimated 1000 p and 1000 pbar). ------------------------------------------------------------
This archive was generated by hypermail 2b29 : Mon Aug 28 2000 - 14:26:21 EDT